OpenStudy (anonymous):

Perform the indicated Operations (a+ 3)^2/ a -3 divided by 5/ (a^2 - 9)

OpenStudy (anonymous):

What i have so far is\[\frac{ (a+3)(a+3) }{ a-3 }\div \frac{ 5 }{ a^2-9 }\]

OpenStudy (anonymous):

\[\frac{(a+3)^2}{\frac{5 (a-3)}{a^2-9}}=\frac{(a+3)^2}{\frac{5 (a-3)}{(a+3) \{a-3\}}}=\frac{1}{5} (a+3)^3 \]a-3 divides out.

OpenStudy (anonymous):

Im still a little confused.

Nnesha (nnesha):

\[\huge\rm \frac{ (a+3)(a+3) }{ a-3 }\div \frac{ 5 }{ a^2-9 }\] change division to multiplication for example \[\huge\rm \frac{ a }{ b } \div \frac{ c }{ d } = \frac{ a }{ b } \times \frac{ d }{ c }\] multiply first fraction by reciprocal of 2nd one

Nnesha (nnesha):

apply difference of square method \[\huge\rm (a^2 - b^2) = (a+b)(a-b)\]